(II) Jamil is purchasing a new truck for $31,500. Jamil is making a $3,000 down payment, and he will make 60 monthly payments of $542 each. What are the total finance costs on this loan?
Answer #1

Cost of truck = $31500

Down payment = $3000

Finance amount = Cost - Down payment

=31500-3000

=28500

Total payment = monthly payment* no of month

=542*60

=32520

Total finance cost = Total payment - Finance amount

=32520-28500

=4020

So finance cost is $4020

Please thumbs up